OpenStudy (anonymous):

solve: x/5+9=4 The solution to 6x – 5 = 6x +2 is x = 8. Select one: True False

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (mrnood):

6x-5=6x+2 implies that -5=2 which is evidently false and has no solution. x/5+9=4 First subtract 9 from each side THEN multiply both sides by 5....

OpenStudy (anonymous):

is it true then

OpenStudy (anonymous):

oh sry i typed the question the rong way